# Parabolic bundles and the intersection cohomology of moduli spaces of vector bundles on curves

### v1

@@ Abstract.
The study of the intersection cohomology of moduli spaces of semistable bundles was initiated by Frances Kirwan in the 1980’s. In this paper, we give a complete geometric proof of a recursive formula, which reduces the calculation of the intersection Betti numbers of the moduli spaces of semistable bundles on Riemann surfaces in degree-0 and arbitrary rank to the known formulas of the Betti numbers of the smooth, degree-1 moduli spaces. Our formula was motivated by the work of Mozgovoy and Reineke in 2015, and appears as a consequence of the Decomposition Theorem applied to the forgetful map from a parabolic moduli space. Our methods are direct, geometric and use a minimal amount of background.
@@ 1. Introduction: the intersection cohomology of moduli spaces
Let $C$ be a smooth complex projective curve of genus $g\geq 2$, and denote by $\mathcal{M}_{d}(r)$ the moduli space of semistable degree-$d$, rank-$r$ vector bundles on $C$. When $r$ and $d$ are coprime, $\mathcal{M}_{d}(r)$ is a smooth projective variety whose cohomology has been computed using a number of techniques, starting in the 1970’s [16, 10, 1, 12, 3]. However, when $\mathrm{gcd}(d,r)>1$, the moduli space $\mathcal{M}_{d}(r)$ is singular, and, in this case, a natural enumerative invariant is intersection cohomology, introduced by Goresky and MacPherson in the late 1970’s [15]. The project to calculate this ”new” cohomology for the moduli spaces $\mathcal{M}_{d}(r)$ of semistable bundles on Riemann surfaces was initiated by Frances Kirwan not long afterwards [18, 20, 19]. One of the most interesting cases is $\mathcal{M}_{0}(r)$ for $r>1$, which is diffeomorphic to the character variety of unitary representations of the fundamental group of $C$ [27]. Over the years, a lot of progress was made, but mostly in the rank-2 case, until in 2015, Mozgovoy and Reineke wrote down a general formula for the intersection Betti numbers of $\mathcal{M}_{d}(r)$ in all ranks and degrees [25]. They used an argument from Donaldson–Thomas theory, which had a gap, however.  [^1] The gap has apparently been fixed recently.
In the present paper, we calculate the intersection Betti numbers of $\mathcal{M}_{0}(r)$ in all ranks using the Decomposition Theorem applied to a particular projection map onto $\mathcal{M}_{0}(r)$, and drawing on some ideas from [32]. We paint a detailed geometric picture of the geometry of parabolic structures on bundles, giving access to Hodge structures, intersection numbers, and other data of the theory. Along the way, we present a fresh look on the Decomposition Theorem itself.
In this introductory chapter, we briefly describe the basic properties of intersection cohomology that we will need, and then give a brisk introduction to the moduli spaces of semistable vector bundles, which is just sufficient to be able to present our main result, Theorem 1.4. This is followed by a description of our strategy, based on a version of the Decomposition Theorem adapted to our needs. For the reader with less patience, the rank-2 case is presented in detail in the following chapter §2.
Acknowledgments The authors wish to thank Ben Davison, Mark De Cataldo, Tamas Hausel, Anton Mellit, Luca Migliorini and Vivek Shende for useful discussions.
@@ 1.2. IC-sheaves
Convention: In this paper, we set the cohomology sheaves of the complex $IC_{Y}$ to be graded from 0 to $2\dim Y$, so that if $Y$ is nonsingular we have $IC_{Y}=\mathbb{Q}_{Y}$.
@@
While in general the intersection cohomology of a variety $Y$ is not easy to determine, we can describe it explicitly when $Y$ is an affine cone over a projective variety (see [23, 11]).
@@
Conventions regarding local systems: Throughout the paper, given a morphism $f:X\rightarrow Y$ of algebraic varieties and an open set $U\subseteq Y$ over which $f$ is a fibration, we will assume that the local systems $R^{j}f_{*}\mathbb{Q}_{f^{-1}(U)}$ are placed in degree $j$ when considered as constructible complexes on $U$. Similarly, we will write $\mathcal{L}^{j}$ for a local system shifted in degree $j$ and, with some abuse of terminology, still refer to a direct sum $\mathcal{L}=\bigoplus_{j}\mathcal{L}^{j}$ as a local system.
@@ 1.3. Semistable bundles
  • As a set $\mathcal{M}_{d}(r)$ is in one-to-one correspondence with the set of (isomoprphism classes) of polystable bundles;
@@ 1.4.1. Notation
    is its Poincaré polynomial, while if $M$ is an algebraic variety, then we introduce
@@ 1.4.2. The plethystic exponential
Consider $\mathbb{Z}^{d}\subset\mathbb{R}^{d}$ as the lattice of possible exponents $\eta$ of monomials $m^{\eta}$, and let $U$ be an open halfspace in $\mathbb{R}^{d}$ containing the origin in its closure. Given a Laurent polynomial (or formal power series) with integer coefficients whose exponent set is supported in $U$
@@ 1.5. The main result
Clearly, formula (5) provides us with a recursion for $IP_{t}(\mathcal{M}_{0}(r))$, if $P_{t}(\mathcal{M}_{1}(r))$ is known.
@@
  • The proof of our formula allows to embed $IH^{\bullet}(\mathcal{M}_{0}(r))$ in $H^{\bullet}(\mathcal{P}_{1}(r))$ a canonical subspace, which gives us direct access to the intersection numbers of $\mathcal{M}_{0}(r)$.
@@ 1.6. The Decomposition Theorem
As mentioned in §1.2, direct calculations of intersection cohomologies are difficult. A more effective strategy is the use of the Decomposition Theorem [2] applied to proper maps $f:X\to Y$, where $X$ is smooth. Traditionally, the Decomposition Theorem is formulated in terms of the semisimplicity of the abelian category of perverse sheaves, and the fact that the simple objects in this category are the intermediate extensions of semisimple local systems.
We will use these results, but instead of investigating the perverse filtration, we will focus on refined methods of calculating the relevant local systems. The Theorem below is a somewhat naive version of the Decomposition Theorem, adapted to our needs. For a more detailed account on the Decomposition Theorem and, broadly, perverse sheaves we refer to De Cataldo and Migliorini [7], who provided a proof via classical Hodge theory in their seminal paper [6], and to the Bourbaki exposé by Williamson [33]
@@
(7) $\mathcal{L}_{\beta}\longrightarrow Y_{\beta}\text{ with fiber }L_{p},\,p\in Y_{\beta}.$
@@
    $L_{p}$ is a graded ideal in $H^{\bullet}(f^{-1}(U))\cong H^{\bullet}(f^{-1}(p))$, whose graded pieces are in the degree range $[\dim X-\dim Y_{\beta}-r(f),\dim X-\dim Y_{\beta}+r(f)]$, where
    $r(f)=\dim X\times_{Y}X-\dim X$ is the defect of semismallness of $f$;
@@
    in the derived category of bounded constructible complexes, there is a (non-canonical) direct sum decomposition
@@
    (1.7.B) $H^{\bullet}(f^{-1}(p))\cong L_{p}\oplus\bigoplus_{Y_{\beta}\in\overline{Y}_{\alpha}}\mathcal{H}^{\bullet}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha})\right)_{p}.$
@@ 1.7. A strategy for calculating $IH(Y)$
    One needs to start with (1.7.B): first, one calculates the cohomologies of the fibers $f^{-1}(p)$ for all $p\in Y$, as well as the stalks of the IC sheaves appearing on the right hand side.
@@
    Armed with these quantities, one can set up a recursive calculation (the subtraction algorithm), which computes all the local system $\mathcal{L}_{\alpha}$ for each stratum $Y_{\alpha}$.
@@
    (8) $IH^{\bullet}(Y,\mathcal{L}_{0})\cong\bigoplus_{i}IH^{\bullet}(Y,\mathbb{Q}_{Y}^{\oplus\dim H^{i}(F_{0})})\cong IH^{\bullet}(Y)\otimes H^{\bullet}(F_{0}),$
    and we need to subtract the contributions of the other strata from the Poincaré polynomial $P_{t}(X)$ in order to isolate (8). In fortunate situations, such as ours, these subtractions lead to another recursion, which, in our case, may be written in the form (5).
@@
  • In §4 we calculate the fiber cohomologies (cf. i. above).
  • In §5, 6,7 we perform the calculation of the local systems (cf. ii.).
  • Finally, in §8, we perform the recursion mentioned in iii. above.
@@ 2. The case of rank 2
While not all complications appear in the $r=2$ case, it is instructive to see our answer explicitly in this case. The final result (10) was first obtained by Kirwan [19].
@@ 2.3. The fibers of the map $\pi$
$\mathcal{P}_{0}(2)=\{(V,F)|\,V-\text{semistable},\deg V=0,\mathrm{rk}V=2,F\subset V_{p}\text{ line},\\ [L\subsetneq V,\deg L=0\Longrightarrow L_{p}\neq F]\}/\cong$
@@
    • similarly, a line bundle over $\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1})$,
@@
    Altogether, this gives us two copies of $\mathbb{P}^{g-1}$ glued together along a single point. Note that the cohomology ring of this space is
@@ 2.5. The subtraction algorithm
$H^{\bullet}(\pi^{-1}(L_{1}\oplus L_{2}))={\mathcal{L}_{[1,1]}}_{L_{1}\oplus L_{2}}\oplus IC(\overline{M}_{[2]},\mathbb{Q})_{L_{1}\oplus L_{2}}.$
@@
of its invariant and anti-invariant parts under the action of $\Sigma_{2}$ exchanging the copies of $\mathcal{M}_{0}(1)$. It follows from the basic properties of intermediate extensions that
@@
Finally, we show that $\mathcal{L}^{\prime}_{[1,1]}=0$, and thus the contribution of $M_{[1,1]}^{\mathrm{nab}}$ to the right hand side of (9) vanishes. We will give a proof of this statement which is valid for general $r$ in §5, but here, for $r=2$ , we present a simple method based on a dimension count.
@@
$i\in[\dim\mathcal{P}_{0}(r)-\dim M_{[1,1]}^{\mathrm{nab}}-r(\pi),\dim\mathcal{P}_{0}(r)-\dim M_{[1,1]}^{\mathrm{nab}}+r(\pi)].$
@@
$\dim\mathcal{P}_{0}(r)=4g-2,\quad\dim M_{[1,1]}^{\mathrm{nab}}=g,\quad\text{and }r(\pi)=1,$
hence the stalk of $\mathcal{L}^{\prime}_{[1,1]}$ at $V$ should lie in $H^{i}(\pi^{-1}(V))$ for $i\in[3g-3,3g-1]$. Since $22\dim\pi^{-1}(V)=2\dim\mathbb{P}^{g-1}=2g-$ 2, this is not possible for $g\geq 2$, hence $\mathcal{L}^{\prime}_{[1,1]}=0$.
@@ 3.1. Stratification of $\mathcal{M}_{0}(r)$
We begin with introducing the necessary notation. A sequence of arbitrary integers will be denoted, as usual, by $(r_{1},\ldots,r_{k})$; we introduce the notation for a partition of the integer $r$:
$\rho=[r_{1},\ldots,r_{k}]\vdash r,$
if $r_{1}+\ldots+r_{k}=r$, and $r_{1}\geq r_{2}\geq\ldots\geq r_{k}>0$. The length of the partition $\rho$, $k$ in this case, will be denoted by $\mathrm{length}(\rho)$.
@@
We have thus $M_{[r]}=\mathcal{M}_{0}(r)^{\mathrm{st}}$. Given two partitions $\mu,\rho\vdash r$, we will write $\mu>\rho$ if $\rho$ is a refinement of $\mu$; in this case, the closure of $M_{\mu}$ contains $M_{\rho}$.

### v2

@@ Abstract.
The study of the intersection cohomology of moduli spaces of semistable bundles was initiated by Frances Kirwan in the 1980’s. In this paper, we give a complete geometric proof of a recursive formula, which reduces the calculation of the intersection Betti numbers of the moduli spaces of semistable bundles on Riemann surfaces in degree-0 and arbitrary rank to the known formulas of the Betti numbers of the smooth, degree-1 moduli spaces. Our formula was motivated by the work of Mozgovoy and Reineke from 2015, and appears as a consequence of the Decomposition Theorem applied to the forgetful map from a parabolic moduli space. We give a detailed description of the topology of this map, and that of the relevant local systems. Our work is self-contained, geometric and focuses on using the multiplicative structures of the theory.
Key words and phrases:
Vector bundles, Parabolic bundles, Intersection cohomology
1991 Mathematics Subject Classification:
14D20
@@ 1. Introduction: the intersection cohomology of moduli spaces
Let $C$ be a smooth complex projective curve of genus $g\geq 2$, and denote by $\mathcal{M}_{d}(r)$ the moduli space of semistable degree-$d$, rank-$r$ vector bundles on $C$. When $r$ and $d$ are coprime, $\mathcal{M}_{d}(r)$ is a smooth projective variety whose cohomology has been computed using a number of techniques, starting in the 1970’s [17, 11, 1, 13, 3]. However, when $\mathrm{gcd}(d,r)>1$, the moduli space $\mathcal{M}_{d}(r)$ is singular, and, in this case, a natural enumerative invariant is intersection cohomology, introduced by Goresky and MacPherson in the late 1970’s [16]. The project to calculate this ”new” cohomology for the moduli spaces $\mathcal{M}_{d}(r)$ of semistable bundles on Riemann surfaces was initiated by Frances Kirwan not long afterwards [20, 22, 21]. One of the most interesting cases is $\mathcal{M}_{0}(r)$ for $r>1$, which is diffeomorphic to the character variety of unitary representations of the fundamental group of $C$ [29]. Over the years, a lot of progress was made, but mostly in the rank-2 case, until in 2015, Mozgovoy and Reineke wrote down a recursive formula for the intersection Betti numbers of $\mathcal{M}_{d}(r)$ in all ranks and degrees [27]. Their work focuses on relating intersection cohomology of stacks to Donaldson–Thomas invariants in the realm of hereditary categories, using Saito’s theory of mixed Hodge modules. Their article was finalized at the end of 2024 while the present work was in the process of being written up. Recently, some of these results were extended to the case of $G$-bundles for an arbitrary reductive group $G$ in [5].
In this paper, we calculate the intersection Betti numbers of $\mathcal{M}_{0}(r)$ in all ranks using the Decomposition Theorem [2] applied to a particular projection map onto $\mathcal{M}_{0}(r)$. Drawing on some ideas from [35], we paint a detailed picture of the geometry of parabolic structures on bundles, giving access to Hodge structures, intersection numbers, and other data of the theory. Our approach is self-contained, mainly uses elementary methods, and gives an explicit description of the local systems appearing in the Decomposition Theorem. We introduce several novel techniques, consistently using multiplicative structures in our work. In particular, this allows us to relate the intersection cohomology of singular moduli spaces to the tautological rings of their smooth counterparts. Along the way, we present a fresh look on the Decomposition Theorem itself.
In this introductory Chapter, we describe the basic properties of intersection cohomology, and then give a brief introduction to the moduli spaces of semistable vector bundles, which is just sufficient to be able to formulate our main result, Theorem 1.4. This is followed by a description of our strategy, based on a version of the Decomposition Theorem adapted to our needs. For the reader with limited patience, we present the rank-2 case in detail in Chapter §2.
Acknowledgments. We are deeply indebted to Donu Arapura and Claude Sabbah for sharing their expertise with us. Useful discussions with Ben Davison, Mark De Cataldo, Tamas Hausel, Anton Mellit, Luca Migliorini, Sergey Mozgovoy, and Vivek Shende are gratefully acknowledged.
@@ 1.2. IC-sheaves
Convention: In this paper, we set the cohomology sheaves of the complex $IC_{Y}$ to be graded from 0 to $2\dim Y$, so that if $Y$ is nonsingular, then $IC_{Y}=\mathbb{Q}_{Y}$.
@@
In general, the intersection cohomology of a variety $Y$ is not easy to determine. When $Y$ is an affine cone over a projective variety $Z$, the following result relates the intersection cohomology of $Y$ to that of $Z$ (see [25, 12]). In particular, when $Z$ is smooth, this provides a viable method for computing the intersection cohomology of $Y$.
@@
Conventions regarding local systems: Throughout the paper, given a morphism $f:X\rightarrow Y$ of algebraic varieties and an open set $U\subseteq Y$ over which $f$ is a fibration, we will assume that the local systems $R^{j}f_{*}\mathbb{Q}_{f^{-1}(U)}$ are placed in degree $j$ when considered as constructible complexes on $U$. Similarly, we will write $\mathcal{L}^{j}$ for a local system shifted in degree $j$ and still refer to a direct sum $\mathcal{L}=\bigoplus_{j}\mathcal{L}^{j}$, with some abuse of terminology, as a local system.
@@ 1.3. Semistable bundles
  • As a set $\mathcal{M}_{d}(r)$ is in one-to-one correspondence with the set of (isomorphism classes) of polystable bundles;
@@ 1.4.1. Notation
    is its Poincaré polynomial, while if $M$ is an algebraic variety, then we introduce the intersection Poincaré polynomial
@@ 1.4.2. The plethystic exponential
Consider $\mathbb{Z}^{d}\subset\mathbb{R}^{d}$ as the lattice of possible exponents $\eta$ of monomials $m^{\eta}$, and let $U$ be an open halfspace in $\mathbb{R}^{d}$ with the origin in its closure. Given a Laurent polynomial (or formal power series) with integer coefficients whose exponent set is supported in $U$
@@ 1.5. The main result
Clearly, formula (5) provides us with a recursion for $IP_{t}(\mathcal{M}_{0}(r))$, if $P_{t}(\mathcal{M}_{1}(r))$ is known [27]
@@
  • The proof of our formula allows us to embed $IH^{\bullet}(\mathcal{M}_{0}(r))$ in $H^{\bullet}(\mathcal{P}_{1}(r))$ as a canonical subspace, which gives us direct access to the intersection numbers of $\mathcal{M}_{0}(r)$.
@@ 1.6. The Decomposition Theorem
Direct calculations of intersection cohomologies are difficult. A more effective strategy is the use of the Decomposition Theorem [2] applied to proper maps $f:X\to Y$, where $X$ is smooth. Traditionally, the Decomposition Theorem is formulated in terms of the semisimplicity of the abelian category of perverse sheaves, and the fact that the simple objects in this category are the intermediate extensions of semisimple local systems.
We will use these results, but instead of investigating the perverse filtration, we will focus on refined methods of calculating the relevant local systems. The Theorem below is a somewhat naive version of the Decomposition Theorem, adapted to our needs. For a more detailed account on the Decomposition Theorem and, broadly, perverse sheaves we refer to De Cataldo and Migliorini [8], who provided a proof via classical Hodge theory in their seminal paper [7], and to the Bourbaki exposé of Williamson [37]
@@
(7) $\mathcal{L}_{\beta}\longrightarrow Y_{\beta}\text{ with stalk }L_{p},\,p\in Y_{\beta}.$
@@
    $L_{p}$ is a graded ideal in $H^{\bullet}(f^{-1}(U))\cong H^{\bullet}(f^{-1}(p))$, whose graded pieces are in the degree range $[\dim X-\dim Y_{\beta}-\boldsymbol{r}(f),\dim X-\dim Y_{\beta}+\boldsymbol{r}(f)]$, where
    $\boldsymbol{r}(f)=\dim X\times_{Y}X-\dim X$ is the defect of semismallness of $f$;
@@
    in the derived category of bounded constructible complexes, there is a direct sum decomposition, depending on the choice of an $f$-ample class,
@@
    (1.7.B) $H^{\bullet}(f^{-1}(p))\cong L_{p}\oplus\bigoplus_{\overline{Y}_{\alpha}\ni p}\mathcal{H}^{\bullet}\left(IC(\overline{Y}_{\alpha},\mathcal{L}_{\alpha})\right)_{p}.$
@@ 1.7. A strategy for calculating $IH(Y)$
    We start with (1.7.B): first, one calculates the cohomologies of the fibers $f^{-1}(p)$ for all $p\in Y$, as well as the stalks of the IC sheaves appearing on the right hand side.
@@
    Armed with these, one can set up a recursive calculation (the subtraction algorithm), which computes the local system $\mathcal{L}_{\alpha}$ for each stratum $Y_{\alpha}$.
@@
    (8) $IH^{\bullet}(Y,\mathcal{L}_{0})\cong IH^{\bullet}(Y)\otimes H^{\bullet}(F_{0}),$
    and we need to subtract the contributions of the other strata from the Poincaré polynomial $P_{t}(X)$ in order to isolate (8). In fortunate situations, these subtractions lead to another recursion, which, in our case, may be written in the form (5).
@@
  • §4 is a technical interlude, describing a Hodge-theoretic refinement of the Decomposition Theorem that we need later.
  • In §5 we calculate the cohomologies of the relevant fibers of our map $\pi$ (cf. i. above).
  • §6, §7, and §8 are devoted to the calculation of the local systems (cf. ii.).
  • Finally, in §9, we construct the recursion mentioned in iii. above.
@@ 2. The case of rank 2
While not all complications appear in the $r=2$ case, it is instructive to see our answer explicitly in this case. The final result (10) was first obtained by Kirwan[21], (see also [19]).
@@ 2.3. The fibers of the map $\pi$
$\mathcal{P}_{0}(2)=\{(V,F)|\,V-\text{semistable},\deg V=0,\mathrm{rk}V=2,F\subset V_{p}\text{ line},\\ \text{ such that }\forall\ L\subsetneq V,\deg L=0\Longrightarrow L_{p}\neq F\}/\cong$
@@
    • a line bundle over $\mathbb{P}\mathrm{Ext}^{1}(L_{2},L_{1})$,
@@
    Altogether, this gives us two copies of $\mathbb{P}^{g-1}$ glued together along a single point. The cohomology ring of this space is
@@ 2.5. The subtraction algorithm
$H^{\bullet}(\pi^{-1}(L_{1}\oplus L_{2}))={\mathcal{L}_{[1,1]}}_{L_{1}\oplus L_{2}}\oplus IC(\overline{M}_{[2]},\mathbb{Q}\oplus\mathbb{Q}[{-2}])_{L_{1}\oplus L_{2}}.$
@@
of its invariant and anti-invariant parts under the $\Sigma_{2}$-action. Then
@@
Finally, we show that $\mathcal{L}^{\prime}_{[1,1]}=0$, and thus the contribution of $M_{[1,1]}^{\mathrm{nab}}$ to the right hand side of (9) vanishes. We will give a proof of this statement which is valid for general $r$ in §6, but here, in the rank-2 case , we present a simple method based on dimension count.
@@
$i\in[\dim\mathcal{P}_{0}(2)-\dim M_{[1,1]}^{\mathrm{nab}}-\boldsymbol{r}(\pi),\dim\mathcal{P}_{0}(2)-\dim M_{[1,1]}^{\mathrm{nab}}+\boldsymbol{r}(\pi)].$
@@
$\dim\mathcal{P}_{0}(2)=4g-2,\quad\dim M_{[1,1]}^{\mathrm{nab}}=g,\quad\text{and }\boldsymbol{r}(\pi)=1,$
hence the stalk of $\mathcal{L}^{\prime}_{[1,1]}$ at $V$ should lie in $H^{i}(\pi^{-1}(V))$ for $i\in[3g-3,3g-1]$. Since $22\dim\pi^{-1}(V)=2\dim\mathbb{P}^{g-1}=2g-$ 2, this is not possible when $g\geq 2$, hence $\mathcal{L}^{\prime}_{[1,1]}=0$.
@@ 3.1. Stratification of $\mathcal{M}_{0}(r)$
We begin with introducing the necessary notation. A sequence of arbitrary integers will be denoted, as usual, by $(r_{1},\ldots,r_{k})$; a partition
$\rho=[r_{1},\ldots,r_{k}]\vdash r$
of the integer $r$ is a decreasing sequence $r_{1}\geq r_{2}\geq\ldots\geq r_{k}>0$ with $r_{1}+\ldots+r_{k}=r$. The length of the partition $\rho$, $k$ in this case, will be denoted by $\mathrm{length}(\rho)$.
@@
In particular, $M_{[r]}=\mathcal{M}_{0}(r)^{\mathrm{st}}$. Given two partitions $\mu,\rho\vdash r$, we will write $\mu>\rho$ if $\rho$ is a refinement of $\mu$; in this case, the closure of $M_{\mu}$ contains $M_{\rho}$.